### v6

# Self-Managing DRAM: A Low-Cost Framework for Enabling Autonomous and Efficient in-DRAM Operations

Abstract
The memory controller is in charge of managing DRAM maintenance operations (e.g., refresh, RowHammer protection, memory scrubbing) to reliably operate modern DRAM chips. Implementing new maintenance operations often necessitates modifications in the DRAM interface, memory controller, and potentially other system components. Such modifications are only possible with a new DRAM standard, which takes a long time to develop, likely leading to slow progress in the adoption of new architectural techniques in DRAM chips.
We propose a new low-cost DRAM architecture, Self-Managing DRAM (SMD), that enables autonomous in-DRAM maintenance operations by transferring the responsibility for controlling maintenance operations from the memory controller to the SMD-chip. To enable autonomous maintenance operations, we tweak the DRAM interface with a single, simple modification, such that an SMD chip rejects memory controller accesses to DRAM regions (e.g., a subarray or a bank) under maintenance, while allowing memory accesses to other DRAM regions. Thus, SMD enables 1) implementing new in-DRAM maintenance mechanisms (or modifying existing ones) with no further changes in the DRAM interface, memory controller, or other system components, and 2) overlapping the latency of a maintenance operation in one DRAM region with the latency of accessing data in another.
We evaluate SMD and show that it 1) can be implemented without modifications to the physical DDRx interface with low latency (0.4% of row activation latency) and area (1.6% of a DRAM chip) overhead, 2) achieves 4.1% average speedup across 20 four-core memory-intensive workloads over a DDR4-based system/DRAM co-design technique that intelligently parallelizes maintenance operations with memory accesses, and 3) guarantees forward progress for rejected memory accesses.
1 Introduction
Advances in manufacturing technology enable increasingly smaller DRAM cell sizes, continuously reducing cost per bit of a DRAM chip [1, 2, 3, 4]. However, as a DRAM cell becomes smaller and the distance between adjacent cells shrink, ensuring reliable and efficient DRAM operation becomes an even more critical challenge [5, 6, 7, 8, 9, 10, 11, 12, 2, 13, 4, 14, 15]. A modern DRAM chip requires three types of maintenance operations (described in detail in section 5) for reliable and secure operation: 1) DRAM refresh [16, 17, 15, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43], 2) RowHammer protection [44, 10, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 40, 56, 57, 58, 59, 60, 61, 62, 63], and 3) memory scrubbing.111Memory scrubbing is not always required in consumer systems but often used in cloud systems [64, 65, 66, 67, 68, 69, 70, 71]. New DRAM chip generations necessitate making existing maintenance operations more aggressive (e.g., lowering the refresh period [72, 73, 44]) and introducing new types of maintenance operations (e.g., targeted refresh [74, 75, 76] and DDR5 RFM [72] as RowHammer defenses).
Two problems likely hinder effective and efficient maintenance mechanisms from being adopted in modern and future DRAM-based computing systems. First, it is difficult to modify existing maintenance mechanisms and introduce new maintenance operations because doing so often necessitates changes to the DRAM interface, which takes a long time (due to various issues related to standardization and agreement across many vendors). Second, it is challenging to keep the overhead of DRAM maintenance mechanisms low as DRAM reliability characteristics worsen and DRAM chips require more aggressive maintenance operations. We expand on the two problems in the next two paragraphs.
Implementing new maintenance operations (or changing existing ones) often necessitates modifications in the DRAM interface, MC, and potentially other system components. Such modifications are only possible with a new DRAM standard, which takes a long time to develop, likely leading to slow progress in the adoption of new architectural techniques in DRAM chips. For example, there was a five-year gap between the DDR3 [77] and DDR4 [78] standards, and eight-year gap between the DDR4 [78] and DDR5 [72] standards. While developing a new standard, DRAM vendors need to push their proposal regarding the maintenance operations through the JEDEC committee (which is the case even if the gap between subsequent DRAM standards were to substantially reduce) such that the new standard includes the desired changes to enable new maintenance operations. Thus, a flexible DRAM interface that decouples improvements to DRAM maintenance operations from interface changes (and hence the timeline needed to enable such changes) would allow DRAM designers to quickly implement custom in-DRAM maintenance mechanisms and develop more reliable and secure DRAM chips.
A maintenance operation triggers more frequently and takes longer with worsening DRAM reliability characteristics and increasing density. While a maintenance operation executes, DRAM is unavailable to serve memory requests (e.g., a periodic refresh operation in DDR4 prevents access to a DRAM rank for up to [78]). As a result, maintenance operations can significantly degrade system performance by delaying critical memory requests. Thus, enabling autonomous maintenance operations without delaying critical memory requests would improve system performance and energy efficiency.
Our goal is to 1) ease and accelerate the process of implementing new in-DRAM maintenance operations and 2) enable more efficient maintenance operations. To this end, we propose Self-Managing DRAM (SMD), a new DRAM architecture that enables implementing new DRAM maintenance operations and modifying the existing ones with a single, simple interface change that eliminates the need for future changes in the DRAM interface, the MC, or other system components.
SMD’s key idea is to allow a DRAM chip to autonomously and efficiently perform maintenance operations by preventing memory accesses only to a relatively small, under-maintenance DRAM region, i.e., a designated section in a DRAM chip (e.g., a DRAM subarray or a bank), while allowing memory accesses to other DRAM regions.222SMD has the key property that the MC does not even know which maintenance operations an SMD chip performs. This property of SMD could also be appealing to DRAM vendors because it would allow vendor-, and chip-generation-specific DRAM resilience characteristics (e.g., the degree and distribution of RowHammer vulnerability and retention failures) to remain undisclosed (i.e., not exposed to other DRAM vendors or system designers).
To prevent access to an under-maintenance DRAM region, an SMD chip rejects a row activation command (i.e., ACT) issued by the MC to the region. To ensure that a DRAM region cannot be locked for too long, SMD enforces a minimum delay (called ARI) between consecutive maintenance operations targeting the same region. We comprehensively study SMD and show that it ensures forward progress for memory requests (section 4.5).
While a DRAM region is under maintenance, the MC can access other regions. This way, a majority of memory accesses are not delayed by maintenance operations, as SMD overlaps the latency of a maintenance operation in a DRAM region with a memory accesses to another DRAM region.
To enable practical adoption of SMD, we implement it with low-cost modifications in existing DRAM chips and MCs. First, to inform the MC that a row activation is rejected, SMD uses a single uni-directional pin on the DRAM chip (which already exists in DDR4/5, see section 7.4). Second, SMD implements a Lock Controller in a DRAM chip for managing regions under maintenance and preventing access to them. Third, SMD adds a new row address latch for every lock region in a DRAM chip to enable accessing one lock region while another is under maintenance (building on the basic design proposed in prior works [79, 80, 16]). In section 7.4, we show that these modifications have low DRAM chip area overhead (1.6% of a DRAM chip) and latency cost (only 0.4% additional row activation latency).
We demonstrate the practicality and versatility of SMD by implementing in-DRAM maintenance mechanisms for DRAM refresh (section 5.1), RowHammer protection (section 5.2), and memory scrubbing (section 5.3). We rigorously evaluate SMD’s impact on system performance and energy efficiency using cycle-accurate memory system simulations (using Ramulator [81, 82]), executing a diverse set of 62 single-core and 60 four-core workloads. We compare SMD-based implementations of the evaluated maintenance mechanisms to their MC-based implementations. We make three key observations from our evaluation. First, an SMD chip that implements lock regions at subarray granularity (i.e., a maintenance mechanism locks a small designated section in a bank) and implements all three maintenance mechanisms (refresh, RowHammer protection, and scrubbing) 1) improves average performance across all memory-intensive four-core workloads by 8.6% compared to a baseline system with a DDR4 chip that supports bank-level refresh and a memory controller that intelligently schedules maintenance operations to avoid delaying main memory accesses (DARP [16]) and 2) outperforms a system in which the DDR4 chip can concurrently perform a maintenance operation in a subarray and a memory access in another subarray (maintenance-access parallelization) and the memory controller can intelligently exploit maintenance-access parallelization (e.g., similar to DSARP’s [16] refresh-access parallelization) by 4.1%.
Second, SMD (with maintenance-access parallelization) reduces DRAM energy consumption by 4.3% on average as it eliminates DRAM commands issued by the memory controller to perform maintenance operations and reduces total execution time. Third, a very low chip area overhead (0.001% of a DRAM chip) implementation of SMD that locks regions at bank granularity (i.e., a maintenance mechanism locks a DRAM bank) and performs only refresh operations for maintenance (similar to DDR4) induces 4.5% average slowdown due to rejected row activations across all memory-intensive four-core workloads compared to the same baseline system (DARP [16]). We conclude that SMD practically provides better performance and energy benefits, without having to modify the DRAM interface (e.g., by adding new timing parameters to the standard) for every new type of maintenance mechanism, than system-DRAM co-design techniques that enable refresh-access parallelization (e.g., [16, 80]).
We expect and hope that SMD will inspire researchers to develop new DRAM maintenance mechanisms that more efficiently tackle the reliability and security problems of DRAM and to enable practical adoption of innovative ideas by providing freedom to both DRAM and MC designers. We will publicly release all SMD source code and data so that others can replicate and build on the work.
We make the following key contributions:
We propose SMD, a new DRAM chip design and interface to enable autonomous and efficient in-DRAM maintenance operations. We implement SMD with small changes to modern DRAM chips and MCs.
We use SMD to implement efficient DRAM maintenance mechanisms for three use cases: DRAM refresh, RowHammer protection, and memory scrubbing.
We rigorously evaluate the performance and energy of SMD-based maintenance mechanisms. SMD provides large performance and energy benefits while also improving reliability across a variety of systems and workloads.
2 Background
2.1 DRAM Organization
A computing system has one or more DRAM channels, where each channel has an independently operating I/O bus. As Fig. 1 illustrates, a Memory Controller (MC) interfaces with one or multiple DRAM ranks via the channel’s I/O bus. A DRAM rank consists of a group of DRAM chips that operate in lockstep. Because the I/O bus is shared across ranks, accesses to different ranks happen in serial manner. A DRAM chip is divided into multiple DRAM banks, each of which is further divided into multiple two dimensional DRAM cell arrays, called subarrays. Within each subarray, DRAM cells are organized as rows and columns. A DRAM row consists of hundreds of DRAM cells which connect to the same wordline. A DRAM cell connects to a sense amplifier via a bitline, and all sense amplifiers in the subarray form a row buffer. Within each DRAM cell, the data is stored as electrical charge on the capacitor and accessed through an access transistor.
2.2 Accessing DRAM
A row must be activated (i.e., opened), to load the row’s data into the row buffer. A row activation is performed by enabling the wordline corresponding to the row address of the ACT command. To enable the corresponding wordline, the global row decoder first partially decodes the row address, and the local row decoder selects a wordline based on the partially decoded address provided by the global row decoder. Enabling the wordline copies the row’s data to the row buffer.
Data in the row buffer is accessible by a RD/WR. The MC follows an ACT with one or multiple RD/WR commands at least after the activation latency (tRCD). The MC issues a PRE command to close a row, after which it can activate a new row from the same bank. The time interval for issuing a PRE after an ACT must be at least restoration latency (tRAS). The MC follows a PRE with an ACT to activate a new row at least after the precharge latency (tRP).
3 Motivation
In current DRAM chips, the MC is in charge of managing DRAM maintenance operations such as periodic refresh, RH protection, and memory scrubbing. When DRAM vendors modify a DRAM maintenance mechanism, the changes often need to be reflected to the MC design as well as the DRAM interface, which makes such modifications very difficult. As a result, implementing new or modifying existing maintenance operations, no matter how fast such implementations or modifications could be developed, can only be realized after a multi-year effort by multiple parties that are part of the JEDEC committee. A prime example to support our argument is the most recent DDR5 standard [72], which took almost a decade to develop after the initial release of DDR4. Even though it might not have taken DRAM designers a decade to develop new maintenance mechanisms, the mechanisms are implemented only after the new standard is released. DDR5 also introduces changes to key issues we study in this paper: DRAM refresh, RH protection, and memory scrubbing. We discuss these changes as motivating examples to show the shortcomings of the status quo in DRAM.
3.1 DRAM Refresh
DDR5 introduces Same Bank Refresh (SBR), which refreshes one bank in each bank group at a time instead of refreshing all banks as in DDR4 [78, 83]. SBR improves bank availability as the MC can access the non-refreshing banks while certain banks are being refreshed. DDR5 implements SBR with a new REFsb command [72]. Implementing REFsb necessitates changes in the DRAM interface and MC.
3.2 RowHammer Protection
In DDR4, DRAM vendors implement in-DRAM RowHammer protection mechanisms by performing Targeted Row Refresh (TRR) within the slack time available when performing regular refresh. However, prior works have shown that in-DRAM TRR is vulnerable to certain memory access patterns [74, 75, 76]. DDR5 specifies the Refresh Management (RFM) that an MC implements to aid in-DRAM RowHammer protection. As part of RFM, the MC uses counters to keep track of row activations to each DRAM bank. When a counter reaches a specified threshold value, the MC issues the new RFM command to DRAM chips. A DRAM chip then performs an undocumented operation with the time allowed for RFM to mitigate RowHammer. Since the first RowHammer work [10] in 2014, it took about 7 years to introduce DDR5 RFM, which requires significant changes in the DRAM interface and the MC design. Still, RFM is likely not a definitive solution for RowHammer as it does not outline a RowHammer defense with security proof. Rather, RFM provides additional time to a DRAM chip for internally mitigating the RowHammer.
3.3 Memory Scrubbing
DDR5 adds support for on-die ECC and in-DRAM scrubbing. A DDR5 chip internally performs ECC encoding and decoding when the chip is accessed. To perform DRAM scrubbing, the MC must periodically issue the new scrub command (for manual scrubbing) or all-bank refresh command (for automatic scrubbing). Similar to Same Bank Refresh and RFM, enabling in-DRAM scrubbing necessitated changes in the DRAM interface and in the MC design.
4 Self-Managing DRAM
To enable autonomously and efficiently performing maintenance operations in a DRAM chip, we introduce the Self-Managing DRAM (SMD) architecture. SMD is part of a continuum between fully master-slave and fully request-reply interfaces. It tries to achieve a balance that enables faster innovation and easier adoption of new ideas.
4.1 Overview of SMD
SMD has a flexible interface that enables efficient implementation of multiple DRAM maintenance operations within the DRAM chip. The key idea of SMD is to provide a DRAM chip with the ability to reject an ACT command via a single-bit negative-acknowledgment (ACT_NACK) signal. An ACT_NACK informs the MC that the DRAM row it tried to activate is under maintenance, and thus temporarily unavailable. Leveraging the ability to reject an ACT, a maintenance operation can be implemented completely within a DRAM chip.
Organization. SMD preserves the general DRAM interface and uses a single uni-directional pin (which already exists in DDR4/5, see section 7.4) in the physical interface of a DRAM chip. This pin is used for transmitting the ACT_NACK signal from the DRAM chip to the MC.
Fig. 2 shows the organization of a bank in an SMD chip. SMD divides a DRAM bank into multiple lock regions of equal size. SMD stores an entry in a per-bank Lock Region Table (LRT) to indicate whether or not a lock region is reserved for performing a maintenance operation.
Operation. A maintenance operation takes place as the Lock Controller sets the lock bit in the LRT entry that corresponds to the lock region on which the maintenance operation is to be performed. When the MC attempts to open a row in a locked region, the Lock Controller generates an ACT_NACK.
A maintenance operation and an ACT can be performed concurrently in the same bank on different lock regions.333An ECC scrubbing operation (section 5.3) uses the data bus, which is a shared resource across all regions. Therefore, a scrubbing operation reservers all lock regions inside a bank at once to prevent any other maintenance operation or memory access from happening at the same time. To enable this, SMD implements a row address latch in order to drive two local row decoders with two independent row addresses.444One row address latch per lock region enables two or more maintenance operations to concurrently happen on different lock regions within a bank. To keep our design simple, we restrict a maintenance operation to happen in one lock region while the MC accesses another lock region. Our design can be easily extended to increase concurrency of maintenance operations across multiple lock regions. When a maintenance operation and an ACT arrive at the same lock region at the same time, the MUX prioritizes the maintenance operation and the SMD chip rejects the ACT.
4.2 Region Locking Mechanism
A maintenance operation can be performed on a lock region only after locking it. Therefore, a maintenance mechanism must lock the region that includes the rows that should undergo a maintenance. A maintenance mechanism can only lock a region that is not already locked, which prevents different maintenance mechanisms from interfering with each other.
Lock Region Size. A lock region consists of a fixed number of consecutively-addressed DRAM rows. To simplify the design, we set the lock region size to one or multiple subarrays. This is because a maintenance operation uses the local row buffers in a subarray, and thus having a subarray shared by multiple lock regions will cause a conflict when accessing a row in a different lock region that maps to the subarray under maintenance. We design and evaluate SMD assuming a default lock region size of 512-row subarrays.
Modern DRAM chips typically use the density-optimized open-bitline architecture [84, 64], which places sense amplifiers on both top and bottom sides of a subarray and adjacent subarrays share sense amplifiers. With the open-bitline architecture, the MC should not access a row in a subarray adjacent to one under maintenance. To achieve this, the Lock Controller simply sends an ACT_NACK when the MC attempts to activate a row in a subarray adjacent to a locked region. Consequently, when SMD locks a region that spans 512-row subarrays, it prevents the MC from accessing subarrays in a bank with rows ( subarrays). In the folded-bitline architecture [84, 64], adjacent subarrays do not share sense amplifiers with each other. Therefore, SMD prevents access only to the subarrays in a locked region. We evaluate SMD using the density-optimized open-bitline architecture.
Ensuring Timely Maintenance Operations. A maintenance mechanism cannot lock a region with an active row. To lock the region, the maintenance mechanism waits for the MC to precharge the row. DRAM standards specify the maximum allowed time for a row to remain active as 9x tREFI [78, 83, 85]). As such, a maintenance mechanism is delayed at most by 9x tREFI.
4.3 Controlling an SMD Chip
While introducing changes in the memory controller (MC) to handle ACT_NACK, SMD also simplifies MC design and operation as the MC no longer implements control logic to periodically issue DRAM maintenance commands. For example, the MC does not 1) prepare a bank for refresh by precharging it, 2) implement timing parameters relevant to refresh, and 3) issue REF commands. The MC still maintains the bank state information (e.g., whether and which row is open in a bank) and respects the DRAM timing parameters associated with commands for performing access (e.g., ACT, PRE, RD, WR).
Handling an ACT_NACK Signal. Fig. 3 depicts a timeline that shows how the MC handles an ACT_NACK signal. Upon receiving an ACT_NACK, the MC waits for ACT Retry Interval (ARI) time to re-issue the same ACT. It keeps re-issuing the ACT command once every ARI until the DRAM chip accepts the ACT. To ensure that a DRAM chip cannot lock a region for prolonged periods, SMD enforces a minimum delay of ARI between the end of a maintenance operation and the start of the next maintenance operation targeting the same region. This allows the MC to serve a request that targets the locked region immediately after the maintenance operation in this region ends. While waiting for ARI, the MC can activate a row from a different lock region or bank, to overlap the ARI latency with a useful operation.
Setting the ACT Retry Interval (ARI). Setting ARI to a very low or high value relative to the expected duration of the maintenance operations can have a negative impact on system performance and energy efficiency. We empirically find that is a favorable configuration for the three maintenance mechanisms that we evaluate in this work (section 5).
Latency. An SMD chip sends an ACT_NACK DRAM command bus cycles after receiving the ACT. should be low so that the MC is quickly notified when an ACT fails, and the MC can attempt to activate a different row in the same bank while waiting for ARI.
The latency has three components: 1) the propagation delay (from the MC to the DRAM chip) of the ACT, 2) the latency of determining whether or not the row to be activated belongs to a locked region, and 3) the propagation delay (from the DRAM chip to the MC) of the signal.
We estimate the overall latency based on the latency of an RD in a conventional DRAM chip. An RD has a latency breakdown that resembles the latency breakdown of a . An RD command 1) propagates from the MC to the DRAM chip, 2) accesses data in a portion of the row buffer in the corresponding bank, and 3) sends the data back to the MC. In the DDR4 standard [78], the latency between issuing an RD and the first data beat appearing on the data bus is defined as tCL (typically 22 cycles for DDR4-3200). The latency components 1) and 3) of RD are similar to those of . Thus, the main difference between and tCL arises from the second component. According to our evaluation, the latency of accessing the Lock Region Table (LRT) is (section 7.4). Given the relatively low complexity of the LRT compared to the datapath that is involved during an RD, we believe the overall latency can be designed to be much smaller than tCL. We assume cycles unless stated otherwise. In our evaluations, we find that small latencies (e.g., ) have negligible effect on system performance mainly because the number of rejected ACTs constitute a small portion of all ACTs.
4.4 ACT_NACK Divergence Across DRAM Chips
SMD maintenance operations take place independently in each DRAM chip. Therefore, when a DRAM rank, which can be composed of multiple DRAM chips operating in lock step, receives an ACT command, some of the chips may send an ACT_NACK while others do not. Normally, ACT_NACK divergence would not happen for maintenance mechanisms that perform the exact same operation at the exact same time in all DRAM chips (e.g., the Fixed Rate Refresh mechanism, Section 5.1). However, a mechanism can also operate differently in each DRAM chip, e.g., a variable rate refresh mechanism [17, 15, 38]. As a result of this divergence, the row becomes partially open in chips that do not send ACT_NACK.
As a solution, the MC waits for ARI and re-issues the same ACT command to retry activating the row in the chips that previously sent an ACT_NACK signal. 555SMD retries activating the DRAM row in chips that rejected the previous activate command by sending a new activate command to all DRAM chips in the rank. DRAM chips that already have an active row simply ignore the activate command. Our analysis shows that under the worse-case ACT_NACK divergence (i.e., when each of the 16 chips in a rank perform refresh on a different region at different time), SMD-VR performs 15.4% worse than SMD-VR with no ACT_NACK divergence, which still provides 7.8% average speedup over conventional DDR4 refresh for memory-intensive four-core workloads.
4.5 Ensuring Forward Progress for Memory Requests
The MC could fail to retry an ACT_NACK’d memory request if an active DRAM row (e.g., in a lock region other than the one targeted by the rejected memory request) precludes retrying the ACT_NACK’d memory request every ARI. Such retry failures could occur repeatedly and lead to temporary starvation of a lock region (i.e., no request targeting that lock region is served for a prohibitively long time). Temporary starvation of a lock region is very unlikely to happen because it requires a maintenance mechanism to repeatedly lock the same lock region for maintenance, which is not a property of a carefully designed maintenance mechanism. Instead, a carefully designed maintenance mechanism alternates between different lock regions every as it performs new maintenance operations (e.g., periodic refresh section 5.1).
To prevent temporary starvation of lock regions irrespective of maintenance mechanism design, SMD issues an ACT_NACK’d memory request every ARI. No memory request is ACT_NACK’d by two separate maintenance operations because an SMD chip does not perform any maintenance operations for ARI after the end of a maintenance operation in a lock region (section 4.3). Issuing a rejected memory request strictly every ARI does not prevent the MC from serving memory requests that access other lock regions. Within one ARI () the MC can 1) activate a DRAM row other than the ACT_NACK’d DRAM row, 2) serve seven (based on the standard values of , , timing parameters) READ/WRITE requests to this newly activated row, and 3) precharge the bank such that an activate command to the previously ACT_NACK’d DRAM row can be issued every ARI.
Proof of Forward Progress. We consider a memory request in the memory controller’s request queue to have made forward progress if it is served by the memory controller.
Theorem. If the MC retries every ARI then makes forward progress.
Proof. We devise a contrapositive proof for the above theorem. Suppose does not make forward progress, i.e., is never dequeued from the request queue. must target a locked region because otherwise is dequeued when it is scheduled. The maintenance mechanism can only hold the lock for finite time. When the maintenance mechanism releases the lock, no maintenance mechanism locks the same region for at least ARI. If the MC retries within one ARI after the lock is released, then is not rejected by a maintenance mechanism, i.e., makes forward progress. However, if the MC retries every ARI, then the MC also retries within one ARI after the lock is released. Thus, if does not make forward progress, then the MC does not retry every ARI. We infer from this that the theorem is true.
Memory Request Stall Time. To limit the maximum potential memory request stall time, 1) long time taking maintenance operations can be divided into many smaller short time taking maintenance operations, 2) the maximum time a maintenance operation takes can be specified in the DRAM standard. In our design, the most time taking maintenance operation is ECC scrubbing (Section 5.3), which takes to read a DRAM row. This stall time is comparable to what a memory request in a modern DDR4-based system experiences. In such a system, a request that arrives at the memory controller’s request queues immediately after the memory controller schedules a periodic refresh operation (by issuing a command) has to wait for the refresh operation to complete (e.g., for [78]).
4.6 Impact to Request Scheduling
The key drawback of our SMD implementation is that it makes the synchronous DDRx interface less predictable. However, the practical performance overheads of this drawback are very low, as we demonstrate in Section 7.
In this paper, we comprehensively describe only one out of many different possible implementations of SMD due to page limits. This implementation prioritizes simplicity. However, the DRAM standard can more strictly specify exactly when and for how long SMD chips are allowed to perform maintenance operations. This stricter implementation would make SMD chips as predictable as DDRx chips today. We hope that future work building on SMD will comprehensively address these issues. To make it more evident that SMD can be implemented in a way that preserves the predictability of the DDRx interface from the perspective of the memory controller (i.e., in a way that keeps the DRAM interface synchronous), we describe a (more) predictable SMD implementation.
Designing a Predictable SMD Interface. A simple way of preserving the DDRx interface’s synchronicity with SMD is to allow the DRAM chip to perform a maintenance operation periodically only at well-defined time intervals. In other words, the DRAM chip can respond with ACT_NACKs only for the fraction of time during which the chip is allowed to execute maintenance operations. We call this fraction of time the maintenance operation time (MOT). Outside of MOT, the DDRx operates as defined today (i.e., synchronously and predictably). During an MOT, the memory controller is allowed to access DRAM cells at the cost of potentially getting ACT_NACK’d by the DRAM chip. The new SMD standard would specify i) how long an MOT is, ii) the period at which the interface enters an MOT, iii) how entry to an MOT is triggered (e.g., by a new MOT command), and iv) how many MOTs the memory controller can postpone.
5 SMD Maintenance Mechanisms
We propose SMD-based maintenance mechanisms for three use cases. However, SMD is not limited to these three use cases and it can be used to support more operations in DRAM.
5.1 Use Case 1: DRAM Refresh
In conventional DRAM, the MC periodically issues REF commands to initiate a DRAM refresh operation. This approach is inefficient due to two main reasons. First, transmitting REFs over the DRAM command bus within the refresh period (e.g., 64, 32, or even depending on the refresh rate) consumes energy and increases the command bus utilization. Due to transmitting a REF over the command bus, the MC may delay a command to another rank in the same channel, which increases DRAM access latency [16]. Second, an entire bank becomes inaccessible while being refreshed although a REF command refreshes only a few rows in the bank. This incurs up to 50% loss in memory throughput [17].
Leveraging SMD, we design a new efficient maintenance mechanism for DRAM refresh: Fixed-Rate Refresh.
Fixed-Rate Refresh (SMD-FR). SMD-FR refreshes DRAM rows in a fixed time interval (i.e., tREFI), similar to conventional DRAM refresh. To limit the time that the MC waits for a region to get unlocked, SMD-FR refreshes (Refresh Granularity) number of rows from a lock region and switches to refreshing another region.
SMD-FR (Fig. 4) operates independently in each DRAM bank and it uses three counters for managing the refresh operations: pending refresh counter, lock region counter, and row address counter. The pending refresh counter is initially zero and SMD-FR increments it by one at the end of a tREFI interval . SMD-FR allows up to refresh operations to be accumulated in the pending refresh counter.666DDR4 [78, 83, 85] allows the MC to postpone issuing up to REF commands in order to serve pending memory requests first. Because the MC can keep a row open for a limited time (section 4.3), the pending refresh counter never exceeds the value .
The lock region and row address counters indicate the next row to refresh. When pending refresh counter is greater than zero, SMD-FR attempts to lock the region indicated by the lock region counter every clock cycle until it successfully locks the region . In some cycles, SMD-FR may fail to lock the region either because the lock region contains an active row or the region is locked by another maintenance mechanism. When SMD-FR successfully locks the region, it initiates a refresh operation that refreshes number of rows in the lock region starting from the row indicated by the row address counter . The refresh operation is conducted by activating and precharging a row to restore the charge in its DRAM cells. We empirically find to be a favorable design point.
After the refresh operation completes, SMD-FR releases the locked region and increments only the lock region counter. When the lock region counter rolls back to zero, SMD-FR also increments by one the row address counter . Finally, SMD-FR decrements the target refresh counter .
Variable Refresh (SMD-VR). In conventional DRAM, all rows are uniformly refreshed with the same refresh period. However, the actual data retention times of different rows in the same DRAM chip greatly vary mainly due to manufacturing process variation and design-induced variation [17, 15, 86, 38, 1, 18]. In fact, only hundreds of “weak” rows across an entire DRAM chip require to be refreshed at the default rate, and a vast majority of the rows can correctly operate when the refresh period is doubled or quadrupled [17]. Eliminating unnecessary refreshes to DRAM rows that do not contain weak cells can significantly mitigate the performance and energy consumption overhead of DRAM refresh [17].
We develop Variable Refresh (SMD-VR), a mechanism that refreshes different rows at different refresh rates depending on the retention time characteristics of the weakest cell in each row. Our SMD-VR design demonstrates the versatility of SMD in supporting different DRAM refresh mechanisms.
The key idea of SMD-VR is to group DRAM rows into multiple retention time bins and refresh a row based on the bin that it belongs to. To achieve low design complexity, SMD-VR uses only two bins: 1) retention-weak rows that have retention time less than and 2) rows that have retention time more than . Inspired by RAIDR [17], SMD-VR stores the addresses of retention-weak rows using a per-bank Bloom Filter [87], which is a space-efficient probabilistic data structure for representing set membership. We assume retention-weak rows are already inserted into the Bloom Filters by the DRAM vendors during post-manufacturing tests.777Alternatively, SMD can be used to develop a maintenance mechanism that performs retention profiling. We leave the development and analysis of it to future work.
The operation of SMD-VR resembles the operation of SMD-FR with the key difference that SMD-VR sometimes skips refreshes to a row that is not in the Bloom Filter, i.e., a row with high retention time. Fig. 5 illustrates how SMD-VR operates. SMD-VR uses the same three counters as SMD-FR. SMD-VR also uses a refresh cycle counter, which is used to indicate the refresh period when all rows (including retention-strong rows) must be refreshed. The refresh cycle counter is initially zero and gets incremented at the end of every refresh period, i.e., when the entire DRAM is refreshed.
SMD-VR increments the pending refresh counter by one at the end of a tREFI interval . When the pending refresh counter is greater than zero, SMD-VR determines whether or not the number of rows, starting from the address indicated by the lock region and row address counters, are retention-weak rows by testing their row addresses using the bank’s Bloom Filter . SMD-VR refreshes the rows that are present in the Bloom Filter every time when it is their turn to be refreshed, as indicated by the lock region and row address counters. In contrast, SMD-VR refreshes the rows that are not present in the Bloom Filter only when the refresh cycle counter has a value that is multiple of , specified as .
Unless stated otherwise, we assume and . Therefore, SMD-VR refreshes the DRAM rows that are not in the Bloom Filter once every four consecutive refresh periods as these rows can retain their data correctly for at least four refresh periods.
After determining which DRAM rows need refresh, SMD-VR operates in a way similar to SMD-FR.
5.2 Use Case 2: RowHammer Protection
Repeatedly activating and precharging (i.e., hammering) a (aggressor) DRAM row causes RowHammer errors in the cells of a nearby (victim) DRAM row [88, 89, 90, 91, 92, 93, 94, 95]. DRAM manufacturers equip their existing DRAM chips with in-DRAM RowHammer protection mechanisms, generally referred to as Target Row Refresh (TRR) [74, 75, 76]. At a high level, TRR protects against RowHammer by detecting an aggressor row and refreshing its victim rows. Because a conventional DRAM chip cannot initiate a refresh operation by itself, TRR refreshes victim rows by taking advantage of the slack time available in the refresh latency (i.e., tRFC), originally used to perform only periodic DRAM refresh [74, 75, 76]. Recent works [74, 75, 76, 96] demonstrate a variety of new RowHammer access patterns that circumvent the TRR protection in chips of all three major DRAM vendors, proving that the existing DRAM interface is not well suited to enable strong RowHammer protection especially as RowHammer becomes a bigger problem with DRAM technology scaling. We use SMD to develop a new maintenance mechanism that overcomes the limitations of the existing TRR mechanisms by initiating victim row refresh within the DRAM chip.
Probabilistic RowHammer Protection (PRP). Inspired by PARA [10], we implement an in-DRAM maintenance mechanism called Probabilistic RowHammer Protection (SMD-PRP). The high level idea is to refresh the nearby rows of an activated row with a small probability. PARA is proposed as a mechanisms in the MC, which makes it difficult to adopt since victim rows are not always known to the MC. SMD enables us to overcome this issue by implementing the PARA-inspired SMD-PRP mechanism completely within the DRAM chip. In addition, SMD-PRP avoids explicit ACT and PRE commands to be sent over the DRAM bus. We omit the implementation details of SMD-PRP due to space limitations.
Fig. 6 illustrates the operation of SMD-PRP.
On a DRAM row activation, SMD-PRP marks the activated row as an aggressor with a small probability of . It marks an aggressor row using a per-bank Marked Rows Table (MRT), that contains an entry for each lock region in the bank. An entry consists of the marked row address and a valid bit. The bit length of the address depend on the size of a lock region. For example, an MRT entry has a 13-bit address field when the lock region size is rows. When MRT contains a marked row, SMD-PRP locks the corresponding region and refreshes the neighbor rows of the marked row . This step can easily accommodate blast radius [54] and any address scrambling. Once the neighbor rows are refreshed, SMD-PRP unlocks the region and unmarks the row in MRT .
SMD-PRP with Aggressor Row Detection (SMD-PRP+). SMD-PRP refreshes victim rows with a small probability on every row activation, even does so for a row that has been activated only a few times. This results in unnecessary victim refreshes, especially for high values that strengthen the RowHammer protection as DRAM cells become more vulnerable to RowHammer with technology scaling.
We propose SMD-PRP+, which detects potential aggressor rows and probabilistically refreshes only their victim rows. The key idea of SMD-PRP+ is to track frequently-activated rows in a DRAM bank and refresh the neighbor rows of these rows using the region locking mechanism of SMD.
SMD-PRP+ tracks frequently-activated rows, within a rolling time window of length , using two Counting Bloom Filters (CBF) [97] that operate in time-interleaved manner. CBF is a Bloom Filter variant that represents the upperbound for the number of times an element is inserted into the CBF. We refer the reader to prior work for background on CBFs [54, 98].
Depicted in Fig. 7, SMD-PRP+ operation is based on 1) detecting a row that has been activated more than times within the most recent interval and 2) refreshing the neighbor rows of this row. must be set according to the minimum hammer count needed to cause a RowHammer bit flip in a given DRAM chip. A recent work shows that activations can cause bit flips in LPDDR4 chips [9]. We conservatively set to . must be equal to or larger than the refresh period in order to track all activations that happen until rows get refreshed by regular refresh operations. We set equal to the refresh period.
Initially, one of the CBFs is in active mode while the other is in passive mode. When activating a row, SMD-PRP+ inserts the address of the activated row to both CBFs . Then, SMD-PRP+ tests the active CBF to check if the accumulated insertion count exceeds the threshold . If so, SMD-PRP+ marks the row in the Marked Rows Table to refresh its neighbor rows with the probability of . SMD-PRP+ does not always mark the row because it is impossible to reset only the counters that correspond to the marked row address in the CBFs. If always marked, a subsequent activation of the same row would again cause the row to be marked, leading to unnecessary neighbor row refresh until all CBF counters are reset, which happens at the end of the interval. After a row is marked, steps - are same as steps - in SMD-PRP. Finally, at the end of an interval, SMD-PRP+ clears the active CBF and swaps the two CBFs to continuously track row activations within the most recent window .
Deterministic RowHammer Protection (DRP). We use SMD to implement SMD-DRP, a deterministic RowHammer protection based on the Graphene [48] mechanism keeps track of frequently activated DRAM rows. Different from Graphene, we implement SMD-DRP completely within a DRAM chip, whereas Graphene requires the MC to issue neighbor row refresh operations.888 We refer the reader to [48] for more details on the Graphene mechanism and its security proof. The operation of SMD-DRP is similar to the operation of Graphene with the difference of SMD-DRP being implemented completely within a DRAM chip, which does not affect the underlying operation, and thus the security proof of Graphene applies to SMD-DRP.
SMD-PRP and SMD-PRP+ are probabilistic mechanisms that provide statistical security guarantees against RowHammer attacks with being the security parameter. On an extremely security-critical system, a deterministic RowHammer protection mechanisms that guarantees the mitigation of RowHammer bit flips at all times may be desired.
The key idea of SMD-DRP is to maintain a per-bank Counter Table (CT) to track the most-frequently activated DRAM rows within a certain time interval (e.g., refresh period of tREFW). Fig. 8 illustrates the operation of SMD-DRP.
When SMD-DRP receives an ACT, it checks if the activated row address () exists in CT . If so, SMD-DRP increments the corresponding CT counter by one . Otherwise, SMD-DRP finds the smallest counter value () in CT and compares it to the value of the spillover counter (SP) , which is initially zero. If SP is equal to , SMD-DRP replaces the row address corresponding to in CT with and increments the corresponding CT counter by one . If SP is smaller than , SMD-DRP increments SP by one . When a CT counter is incremented in step , SMD-DRP checks if the counter value is a multiple of , which is the maximum number of times a row can be activated without refreshing its neighbors. If so, SMD-DRP refreshes the neighbors of . To prevent the counters from overflowing, SMD-DRP resets the CT counters and SP on every tREFW interval.
To ensure that no row is activated more than without refreshing its neighbor rows, the number of CT counters () must be configured as follows:
where is the maximum number of activations that the MC can perform within a tREFW interval in a single bank. In our evaluations, we set to show that the overheads of SMD-DRP are small even for DRAM chips that are extremely vulnerable against RowHammer.
5.3 Use Case 3: Memory Scrubbing
To mitigate the increasing bit errors mainly caused by the continued DRAM technology scaling, DRAM vendors equip their DRAM chips with on-die Error Correction Codes (ECC) [99, 100, 101, 8]. On-die ECC is designed to correct a single bit error assuming that a failure mechanism is unlikely to incur more than one bit error in a codeword [100]. However, even when the assumption always holds, a failure mechanism can gradually incur two or more bit errors over time. A widely-used technique for preventing the accumulation of an uncorrectable number of bit errors is memory scrubbing. Memory scrubbing describes the process of periodically scanning the memory for bit errors in order to correct them before more errors occur.
We propose SMD-based Memory Scrubbing (SMD-MS), which is an in-DRAM maintenance mechanism that periodically performs scrubbing on DRAM chips with on-die ECC.999To preserve ECC transparency and enable system-level policies that depend on ECC scrubbing information (e.g., detecting and replacing faulty DRAM chips), SMD-MS exposes critical ECC scrubbing information (e.g., the address of the DRAM row with the most corrected errors) to the memory controller via DRAM mode-status registers, similar to how ECC Error Check and Scrub does in DDR5 [72]. Compared to conventional MC-based scrubbing, SMD-MS eliminates moving data to the MC by scrubbing within the DRAM chip, and thus reduces the performance and energy overheads of memory scrubbing.
SMD-MS operation resembles the operation of SMD-FR. Similar to SMD-FR, SMD-MS maintains pending scrub counter, lock region counter, and row address counter. SMD-MS increments the pending scrub counter at fixed intervals of . When the pending scrub counter is greater than zero, SMD-MS attempts to lock the region indicated by the lock region counter. After locking the region, SMD-MS performs scrubbing operation on the row indicated by the row address counter. The scrubbing operation takes more time than a refresh operation as performing scrubbing on a row consists of three time consuming steps for each codeword in the row: 1) reading the codeword, 2) performing ECC decoding and checking for bit errors, and 3) encoding and writing back the new codeword into the row only when the decoded codeword contains a bit error. Refreshing a row takes , whereas scrubbing a row takes when no bit errors are detected.101010We assume ECC decoding/encoding hardware is fully pipelined. In case of a bit error, writing a corrected codeword incurs latency. Therefore, SMD-MS keeps a region locked for longer than SMD-FR. When the scrubbing operation is complete, SMD-MS releases the lock region and increments the lock region and row address counters, and decrements the pending scrub counter as in SMD-FR.
Rank-Level Memory Scrubbing. Server memory modules typically use rank-level ECC [102, 103, 104]. We discuss why rank-level ECC and in-DRAM ECC scrubbing should be done separately but in a combined way in server modules. During a conventional DRAM RD operation, a DRAM chip with in-DRAM ECC (e.g., a DDR5 chip [72]) performs error correction for the in-DRAM ECC codeword without 1) writing the corrected codeword back and 2) keeping a record of the correction in mode-status registers [72]. Thus, in a rank-level ECC scrubbing operation, the memory controller cannot distinguish between an in-DRAM-ECC-corrected codeword and a codeword with no errors.
In a computing system that uses memory modules with in-DRAM ECC, relying only on the rank-level ECC scrubbing to prevent the accumulation of errors is costly. This is because the memory controller must write back every codeword that it reads from DRAM for every rank-level ECC scrubbing operation. Doing so could incur performance overheads due to the long data bus turn-around latency in DRAM (e.g., in DDR5 is approximately [105, 72]) and energy overheads due to increased data movement between the memory controller and the DRAM chips. Instead, using SMD, the computing system could use both rank-level ECC scrubbing and in-DRAM ECC scrubbing to 1) leverage the (typically) stronger error correction capability of rank-level ECC and 2) leverage the performance and energy efficiency of in-DRAM ECC while preventing the accumulation of errors.
5.4 Other Use Cases
To demonstrate that SMD is not limited to the three use cases we primarily focus on, we discuss two other use cases.
Online DRAM Error Profiling. SMD enables implementing a maintenance mechanism for online profiling of DRAM errors that may occur when operating with increased refresh period and reduced DRAM timing parameters [34, 17, 106, 107, 108, 1, 109, 110, 86, 38, 15, 111]. Such a maintenance mechanism can be used to explore how DRAM cell characteristics vary throughout the DRAM chip. To profile a DRAM region, the maintenance mechanism can use SMD to lock this region and prevent the MC from accessing it while the region is being profiled. Because the profiling operation can cause bit flips, the maintenance mechanism should temporarily buffer the original data that is stored in the profiled region. For example, for profiling errors that occur at reduced tRCD, the maintenance mechanism should first safely copy the data stored in a row to another storage space (e.g., an unused row or an SRAM buffer), before attempting to activate the row with reduced tRCD. The profiler mechanism can internally share the profiling results with other maintenance mechanisms (e.g., for skipping refreshes to rows with high retention time) and mechanisms for improving DRAM access latency and energy efficiency.
Processing in/near Memory. In the presence of an in-DRAM processing engine, SMD can help resolve access conflicts between the in-DRAM processing engine and the MC. To do so, SMD can treat the in-DRAM processing engine as a maintenance mechanism. The in-DRAM processing engine can use SMD to lock a DRAM region that it will operate on. Because SMD does not allow the MC to activate a row in a locked region, only the in-DRAM processing engine will have access to the locked region until it completes the processing and releases the region.
6 Experimental Methodology
We extend Ramulator [82, 81] to implement and evaluate the three SMD maintenance mechanisms (SMD-FR, SMD-DRP, and SMD-MS) that we describe in section 5. We use DRAMPower [112, 113] to evaluate DRAM energy consumption. We use Ramulator in CPU-trace driven mode executing traces of representative sections of our workloads collected with a custom Pintool [114]. We warm-up the caches by fast-forwarding 100 million (M) instructions. We simulate each representative trace for 500M instructions (for multi-core simulations, until each core executes at least 500M instructions).
We use the system configuration provided in Table 1 in our evaluations. Although our evaluation is based on DDR4 DRAM, the modifications required to enable SMD can be adopted in other DRAM standards, as we explain in section 7.4.
Processor | & 4-wide issue CPU core, 1-4 cores, 8 MSHRs/core, 128-entry instruction window |
Last-Level Cache | cache-line, 8-way associative, /core |
Memory Controller | 64-entry read/write request queue, FR-FCFS-Cap [115] |
DRAM | DDR4-3200 [78], refresh period, 4 channels, 2 ranks, 4/4 bank groups/banks, 128K-row bank, 512-row subarray, row size |
DDR4 Baseline. The baseline system 1) uses per-rank refresh111111The DDR4 standard does not support per-bank refresh. and 2) does not perform ECC-scrubbing.
Workloads. We evaluate 62 single-core applications from four benchmark suites: SPEC CPU2006 [116], SPEC CPU2017 [117], TPC [118], STREAM [119], and MediaBench [120]. We classify the workloads in three memory intensity groups measured using misses-per-kilo-instructions (MPKI) in the last-level cache (LLC): low (), medium (), and high (). We randomly combine single-core workloads to create multi-programmed workloads. Each multi-programmed workload group, 4c-low, 4c-medium, and 4c-high, contains four-core workloads.
Metrics. We use Instructions Per Cycle (IPC) to evaluate the performance of single-core workloads. For multi-core workloads, we evaluate the system throughput using the weighted speedup metric [121, 122, 123].
Comparison Points. We compare SMD to memory controller/DRAM co-design techniques. First, DARP [16] intelligently schedules per-bank refresh commands to idle banks while the memory controller is in write mode to reduce delays imposed by refresh operations on memory demand requests (e.g., load instructions executed by the processor). Second, DSARP [16] implements DARP and modifies DRAM chips and the interface to perform a refresh operation in one subarray and a memory accesses in another subarray and thereby hide the latency of refresh operations and reduce delays they impose on memory requests. We modify DARP and DSARP such that they consider victim row refresh operations issued by Deterministic RowHammer Protection (section 5.2) and attempt to reduce delays these operations impose on memory requests.
Configuring Maintenance Mechanisms. SMD-FR refreshes a DRAM row every . Based on [17], for SMD-VR, we conservatively assume 0.1% of rows in each bank need to be refreshed every while the rest retain their data correctly for and more. SMD-VR uses a -bit Bloom Filter with hash functions. SMD-PRP and SMD-PRP+ refresh the victims of an activated row with a high probability, i.e., . SMD-DRP refreshes the victims before the aggressor is activated (512) times during a tREFW (). SMD-MS operates with an aggressive 5-minute scrubbing period.
Evaluated System Configurations. SMD-FR-1LR implements a single lock region in a bank, i.e., it cannot concurrently perform a refresh operation and a memory access. DSARP/SMD-FR can concurrently perform a refresh in a subarray/lock region and a memory access in another. MC-Combined-DARP, MC-Combined-DSARP, and SMD-Combined implement the memory-controller-(MC) and SMD-based versions of Fixed Rate Refresh (section 5.1), Deterministic RowHammer Protection (section 5.2), and Memory Scrubbing (section 5.3). MC-based implementations of SMD maintenance mechanisms are functionally equivalent to SMD-based ones (e.g., SMD-DRP and MC-DRP are configured to provide the same RowHammer prevention guarantees).
7 Evaluation
We evaluate the performance and energy efficiency of SMD-based maintenance mechanisms.
7.1 Single-core Performance
Fig. 9 shows the speedup of the 22 highest MPKI single-core workloads for DARP, DSARP, SMD, and a hypothetical No-Refresh configuration that does not perform any maintenance operations, over the Baseline system.
We make three key observations. First, SMD-Combined provides a comparable average speedup of 5.9% to No-Refresh’s 5.0% as SMD-Combined eliminates a large fraction of the maintenance overhead by allowing the memory controller to access lock regions that are not under maintenance. Second, SMD-Combined outperforms MC-Combined-DSARP and MC-Combined-DARP by 0.6% and 3.9%, respectively. While MC-Combined-DSARP can concurrently access main memory and perform a maintenance operation, it needs to issue a DRAM command (e.g., a per bank refresh command for periodic refresh) for each maintenance operation and incur delays for other DRAM commands that the memory controller issues to serve main memory requests. In contrast, SMD-Combined autonomously performs each maintenance operation inside the DRAM chip and does not incur delays for DRAM commands (except those incurred by ACT_NACKs). MC-Combined-DARP cannot concurrently perform a maintenance operation and access main memory, thereby performs worse than both SMD-Combined and MC-Combined-DSARP. We make similar observations on SMD-FR’s, DSARP’s, and DARP’s performance. These observations are in line with prior work [16]. Third, intelligent refresh scheduling (DARP) provides 1.1% average speedup across all workloads by reducing the impact of refresh operations on the latency of memory requests. SMD-FR-1LR incurs 2.9% slowdown over DARP. Although SMD-FR-1LR locks a bank only when there is no active row in a bank, it does not have a global view over memory requests in memory controller read/write queues, unlike DARP, and cannot opportunistically perform refresh operations while the memory controller is in write mode.
7.2 Multi-core Performance
Fig. 10 shows weighted speedup (normalized to the weighted speedup of the DDR4 baseline) for 60 four-core workloads (20 per memory intensity level). The white circles inside each box (error lines) represent the average (minimum and maximum) weighted speedup across the 20 workloads in the corresponding group.
We make three major observations from Fig. 10. First, SMD-Combined provides 8.6% and 4.1% speedups on average across 4c-high workloads over MC-Combined-DARP and MC-Combined-DSARP, respectively. We attribute these speedups to i) SMD’s maintenance-access parallelization and ii) SMD’s ability to perform maintenance operations autonomously inside the DRAM chip without the memory controller having to issue DRAM commands. Second, SMD-FR-1LR incurs an average 4.5% slowdown on average across 4c-high workloads compared to DARP. While DARP can reduce the overheads of periodic refresh by intelligently scheduling refresh operations, SMD-FR-1LR cannot. We attribute SMD-FR-1LR’s overheads to the relatively high rate of ACT_NACK commands (not shown in the figure) it issues: SMD-FR-1LR issues an ACT_NACK command for every 11.7 activate commands, on average across all 4c-high workloads. Third, we observe that i) the speedup and slowdown trends for the tested mechanisms are similar for the average 4c-medium workload, and ii) the tested mechanisms provide little performance improvement for the average 4c-low workload with very low memory intensity.
We conclude that SMD i) provides substantial system performance benefits by concurrently performing a maintenance operation and a memory access in different lock regions in a bank and ii) does not significantly hurt system performance when SMD performs maintenance at bank granularity (i.e., implements one lock region per DRAM bank).
7.3 SMD’s Energy Consumption
Fig. 11 shows the average DRAM energy consumption normalized to the DDR4 baseline for single- and four-core workloads. We make two major observations.
First, SMD-Combined and SMD-FR reduce DRAM energy by 3.9% (4.3%) and 5.2% (5.2%) across all single-core (4c-high) workloads compared to the baseline, respectively. SMD-Combined performs close to the hypothetical No-Refresh configuration, providing 59.7% of the energy reduction benefits of No-Refresh on average across 4c-high workloads. We attribute the reduction in DRAM energy to i) the reduced DRAM background energy consumption because SMD shortens the execution time for these workloads and ii) eliminating DRAM commands for maintenance operations (e.g., for periodic refresh) on the power-hungry DDRx bus. Second, SMD-FR-1LR increases average single-core (4c-high) workload DRAM energy consumption by 0.1% (3.0%). While SMD-FR-1LR also eliminates DRAM commands for maintenance operations, it induces energy overheads mainly because of the increased execution time for the evaluated workloads.
7.4 Hardware Overhead
An SMD chip needs to transmit ACT_NACK signals to the memory controller (MC). SMD can 1) repurpose the existing alert_n pin in DDR4 chips [78, 124] or 2) introduce an extra physical pin to transmit ACT_NACK signals.
1) alert_n. alert_n is currently used to inform the MC that the DRAM chip detected a CRC or parity check failure on the issued command, and thus the MC must issue the command again. The alert_n signal can simply be asserted not only on CRC or parity check failure, but also when the MC attempts accessing a row in a locked region. This approach does not require an additional physical pin. However, 1) assigning multiple meanings to an already ambiguous alert_n signal [78] could complicate MC design and 2) alert_n, as currently defined in the standard [78], is an open drain signal that is slow and not precise enough.
2) Introducing a new pin. To potentially simplify MC design, SMD can introduce a new pin to transmit ACT_NACK signals. Only one new pin between the MC and all memory devices (e.g., all DRAM chips in a memory channel or a rank) is enough for systems that use rank-based DRAM chip organizations (e.g., DDR4 DIMMs). For example, 128 new pins are needed for a high-end system with 32 memory channels equipped with 4-rank memory chips in each channel. However, a system of this scale already has 4’094 processor pins [125], and 128 new pins would amount to a relatively small 3.13% increase in pins. Similar to how alert_n signal works, the module can send the MC a single ACT_NACK when any of the per-chip ACT_NACK signals are asserted.
We individually discuss the other changes required on the existing DRAM chip and the MC circuitry.
DRAM Chip Modifications. We use CACTI [126] to evaluate the hardware overhead of the changes that SMD introduces over a conventional DRAM bank (highlighted in Fig. 2) assuming technology.121212Capitalizing on the latest DRAM technology libraries to implement the changes that SMD introduces would likely provide more accurate area overhead estimations. However, such libraries and tools are proprietary and we do not have access to them. Lock Region Table (LRT) is a small table that stores a single bit for each lock region in a DRAM bank to indicate whether or not the lock region is under maintenance. In our evaluation, we assume that a bank is divided into lock regions. Therefore, LRT consists of only bits, which are indexed using a 4-bit lock region address. According to our evaluation, an LRT incurs only area overhead per bank. The area overhead of all LRTs in a DRAM chip is only 0.001% of a DRAM chip. The access time of an LRT is , which is only 0.4% of typical row activation latency (tRCD) of .
SMD adds a per-region RA-latch to enable accessing one lock region while another is under maintenance. An RA-latch stores a pre-decoded row address, which is provided by the global row address decoder, and drives the local row decoders where the row address is fully decoded. This design builds on the basic design proposed in SALP [79] and refresh-access parallelization introduced in [16, 80]. According to our evaluation, all RA-latches incur a total area overhead of 1.6% of a DRAM chip. An RA-latch has only latency, which is negligible compared to tRCD.
Besides these changes that are the core of the SMD substrate, a particular maintenance mechanism may incur additional area overhead. We evaluate the DRAM chip area overhead of the maintenance mechanisms presented in section 5. The simple refresh mechanism, SMD-FR, requires only additional area in a DRAM chip. The DRAM chip area overhead of the refresh mechanism is less than 0.1% of a typical DRAM chip. SMD-DRP requires a large Counter Table with 1224 counters per bank for the RowHammer threshold value that we use in our performance evaluation. Across a DRAM chip, SMD-DRP requires area, which is 7.0% of a typical DRAM chip size. The control logic of SMD-MS is similar to the control logic of SMD-FR and it requires only additional area excluding the area of the ECC engine, which is already implemented by DRAM chips that support in-DRAM ECC.
Memory Controller Modifications. We slightly modify the MC’s scheduling mechanism to retry a rejected ACT command as we explain in section 4.3. Upon receiving an ACT_NACK, the MC marks the bank as precharged. An existing MC already implements control circuitry to pick an appropriate request from the request queue and issue the necessary DRAM command based on the DRAM bank state (e.g., ACT to a precharged bank or RD if the corresponding row is already open) by respecting the DRAM timing parameters. The ACT Retry Interval (ARI) is simply a new timing parameter that specifies the minimum time interval for issuing an ACT to a lock region after receiving an ACT_NACK from the same region. Therefore, SMD can be implemented in existing MCs with only slight modifications by leveraging the existing request scheduler (e.g., FRFCFS [115]).
Tracking locked regions. To apply the ARI timing parameter (e.g., to an ACT command targeting a locked region), the memory controller tracks which regions are locked. The memory controller could store the address of the locked region in every DRAM bank (since only one region can be under maintenance at a time in a bank). 1 rank address bit, 4 bank address bits, and 4 lock region address bits (9 bits) are sufficient to track every under-maintenance lock region in 2 ranks, 8 chips, and 16 banks (256 unique banks). Therefore, the storage cost for tracking locked regions is only 288 bytes for the evaluated memory channel with a dual-rank x8 memory module.
Address mapping schemes. SMD does not require modifications to how a physical address is mapped to a memory-controller-visible DRAM address. In section 7, we evaluate SMD using a mapping scheme that interleaves consecutive cache blocks across DRAM channels (from the most significant to the least significant bit, the physical address is mapped to rows, banks, ranks, columns, and channels). We also evaluate SMD using the mapping scheme described in [127] that aims to exploit bank-level parallelism. For this mapping scheme, SMD-FR provides 8.7% speedup over the baseline system, on average across 4c-high workloads. We leave detailed evaluation of SMD with different address mapping schemes for future work.
No further changes are required in the MC to support different maintenance mechanisms enabled by SMD. Thus, SMD enables DRAM designers to update existing maintenance mechanisms and implement new ones without any further changes to MCs that support SMD.
7.5 Sensitivity Studies
We analyze the performance and DRAM energy effects of different SMD configuration parameters.
7.5.1 DRAM Refresh Period
Fig. 12 plots the speedup and energy reduction that SMD-FR, SMD-VR, and SMD-combined achieve over DDR4 for different refresh periods across single- and four-core workloads. We make three key observations. First, the performance and energy benefits of SMD-FR/SMD-VR increase as the refresh period reduces, achieving 50.5%/53.6% speedup and 25.6%/29.0% DRAM energy reduction on average at refresh period across 4c-high workloads. Thus, SMD-based refresh mechanisms will become even more beneficial to employ in future DRAM chips that are expected to require more frequent refresh [17, 8, 86]. Second, SMD-combined provides 51.5% speedup and 27.9% DRAM energy reduction on average across 4c-high at refresh period, showing that the overheads of SMD-based RowHammer protection and memory scrubbing mechanisms remain low even at high refresh rates. Third, all SMD-based maintenance mechanism eliminate most of the performance overhead of refresh across all refresh periods (e.g., SMD-combined reaches 95.6% of the hypothetical NoRefresh DRAM at ).
We conclude that SMD-based maintenance mechanisms improve DRAM reliability and security while greatly reducing the overhead of DDR4 refresh. Their performance benefits are expected to increase for future DRAM chips that are likely to require high refresh rates.
7.5.2 ACT_NACK Divergence Across Chips
In Section 4.4, we explain divergence in SMD maintenance operations can happen when different DRAM chips in the same rank perform maintenance operation at different times. Such a divergence leads to a partial row activation when the activated row is in a locked region in some DRAM chips but not in others. To handle partial row activations, we develop three policies.
Precharge. With the Precharge policy, the MC issues a PRE command to close the partially activated row when some DRAM chips send ACT_NACK but others do not. After closing the partially activated row, the MC can attempt to activate a row from a different lock region in the same bank.
Wait. With the Wait policy, the MC issues multiple ACT commands until a partially activated row becomes fully activated. When some chips send ACT_NACK for a particular ACT but others do not, the MC waits for ARI and issues a new ACT to attempt activating the same row in DRAM chips that previously sent ACT_NACK.
Hybrid. We also design a Hybrid policy, where the memory controller uses the Precharge policy to close a partially activated row if the request queue contains or more requests that need to access rows in different lock regions in the same bank. If the requests queue has less than requests to different lock regions, the MC uses the Wait policy to retry activating the rest of the partially activated row.
In Fig. 13, we compare the performance and energy savings of SMD-FR and SMD-VR when using the three ACT_NACK divergence handling policies across 4c-high workloads. The plots show results for the common-case (CC) and worst-case (WC) scenarios with regard to when maintenance operations happen across different SMD chips in the same rank. In the common-case scenario, the DRAM chips generally refresh the same row at the same time due to sharing the same DRAM architecture design. However, refresh operations in some of the DRAM chips may still diverge during operation depending on the refresh mechanism that is in use. For example, SMD-VR refreshes retention-weak rows, whose locations may differ across the DRAM chips, at a higher rate compared to other rows, resulting in divergence in refresh operations across the DRAM chips in a rank. In the worst-case scenario, we deliberately configure the DRAM chips to refresh different rows at different times. For this, we 1) delay the first refresh operation in by , where and is the latency of a single refresh operation, and 2) set the Lock Region Counter (LRC) of to .
We make three key observations. First, the performance and DRAM energy consumption only slightly varies across different divergence handling policies in both common-case and worst-case scenarios. This is because, after a partial row activation happens, a different row that is not in a locked region in all off of the DRAM chips often does not exist in the memory request queue. As a result, the Precharge and Hybrid policies perform similarly to the Wait policy.
Second, SMD-FR performs worse than the DDR4 baseline in the worst-case refresh distribution scenario (i.e., average slowdown of 2.6% with the Wait policy). Certain individual workloads experience even higher slowdown (up to 12.8% with the Wait policy). The reason is that, when different DRAM chips lock a region at different times, the total duration during which the lock region is unavailable becomes times the duration when all chips simultaneously refresh the same lock region. This significantly increases the performance overhead of refresh operations.
Third, SMD-VR does not suffer much from the divergence problem and outperforms the DDR4 baseline even in the worst-case scenario. This is because SMD-VR mitigates the DRAM refresh overhead by significantly reducing the number of total refresh operations by exploiting retention-time variation across DRAM rows. Thus, the benefits of eliminating many unnecessary refresh operations surpass the overhead of ACT_NACK divergence.
We conclude that 1) although SMD-FR suffers from noticeable slowdown for the worst-case scenario, which should not occur in a well-designed system, it still provides comparable performance to conventional DDR4 and 2) SMD-VR outperforms the baseline and saves DRAM energy even in the worst-case scenario.
7.5.3 Comparison to Conventional Scrubbing
Fig. 14 compares the performance and energy overheads of conventional DDR4 scrubbing to SMD-MS across 4c-high workloads. In the figure, DDR4 Scrubbing represents the performance and energy overhead of conventional scrubbing compared to DDR4 without memory scrubbing. Similarly, SMD-MS represents the performance and DRAM energy overhead of SMD-based scrubbing compared to SMD-FR.
We make two observations. First, both DDR4 scrubbing and SMD-MS have negligible performance overhead for scrubbing periods of 5 minutes and larger because scrubbing operations are infrequent at such periods. Second, DDR4 scrubbing causes up to 1.49%/8.84% average slowdown for 1 minute/10 second scrubbing period (up to 2.37%/14.26%), while SMD-MS causes only up to 0.34%/1.78% slowdown. DDR4 scrubbing has high overhead at low scrubbing periods because moving data from DRAM to the MC to perform scrubbing is inefficient compared to performing scrubbing within DRAM using SMD-MS. Scrubbing at high rates may become necessary for future DRAM chips as their reliability characteristics continuously worsen. Additionally, mechanisms that improve DRAM performance at the cost of reduced reliability [15, 128] can use frequent DRAM scrubbing to achieve the desired DRAM reliability level.
We conclude that SMS performs memory scrubbing more efficiently than conventional MC based scrubbing and it enables scrubbing at high rates with small performance and energy overheads.
7.5.4 Comparison to PARA in Memory Controller
Fig. 15 compares the performance and energy overheads of PARA implemented in the MC (as proposed by Kim et al. [10]) for DDR4 and SMD-PRP for different neighbor row activation probabilities (i.e., ) across 4c-high workloads. PARA represents the performance and energy overheads with respect to conventional DDR4 with no RowHammer protection. Similarly, SMD-PRP represents the performance and energy overheads with respect to a SMD chip, which uses SMD-FR for periodic refresh, with no RowHammer protection.
We make two observations. First, the performance and energy consumption of MC-based PARA scales poorly with . At the default of 1%, PARA incurs 12.4%/11.5% average performance/DRAM energy overhead. For higher , the overheads of PARA increase dramatically to 81.4%/368.1% at of 20%. Second, SMD-PRP is significantly more efficient than PARA. At the default of 1%, SMD-PRP incurs only 1.4%/0.9% performance/DRAM energy overheads and at of 20% the overheads become only 11.3%/8.0%. SMD-PRP is more efficient than PARA mainly due to enabling access to non-locked regions in a bank while SMD-PRP performs neighbor row refreshes on the locked region.
We conclude that SMD-PRP is a highly-efficient RowHammer protection that incurs small performance and DRAM energy overheads even with high neighbor row refresh probability, which is critical to protect future DRAM chips that may have extremely high RowHammer vulnerability.
7.5.5 SMD-DRP Maximum Activation Threshold
We analyze the SMD-DRP’s sensitivity to the maximum row activation threshold (). Fig. 16 shows the average speedup that SMD-FR and SMD-DRP achieve for different values across 4c-high workloads compared to the DDR4 baseline. When evaluating SMD-DRP, we use SMD-FR as a DRAM refresh mechanism.
We observe that SMD-DRP incurs negligible performance overhead on top of SMD-FR even for extremely small values. This is because SMD-DRP generates very few neighbor row refreshes as 4c-high is a set of benign workloads that do not repeatedly activate a single row many times.
Although the performance overhead of SMD-DRP is negligible, the number of Counter Table (CT) entries required is significantly large for small values. For , SMD-DRP requires counters per bank, and the number of counters required increase linearly as reduces, reaching counters at the lowest that we evaluate.
7.5.6 Number of Vulnerable Neighbor Rows
We analyze the performance overheads of SMD-PRP and SMD-DRP when refreshing a different number of neighbor rows upon detecting a potentially aggressor row. Kim et al. [9] show that, in some DRAM chips, an aggressor row can cause bit flips also in rows that are at a greater distance than the two victim rows surrounding the aggressor row. Thus, it may be desirable to configure a RowHammer protection mechanism to refresh more neighbor rows than the two rows that are immediately adjacent to the aggressor row.
Fig. 17 shows the average speedup that SMD-Combined (separately with SMD-PRP and SMD-DRP) achieves for different number of neighbor rows refreshed across 4c-high workloads compared to the DDR4 baseline. The Neighbor Row Distance values on the x-axis represent the number of rows refreshes on each of the two sides of an aggressor row (e.g., for neighbor row distance of , SMD-PRP and SMD-DRP refresh four victim rows in total).
We make two key observations from the figure. First, SMD-PRP incurs large performance overheads as the neighbor row distance increases. This is because, with , SMD-PRP perform neighbor row refresh approximately for every 100th ACT command, and the latency of this refresh operation increases with the increase in the number of victim rows. Second, the performance overhead of SMD-DRP is negligible even when the neighbor row distance is five. This is because, SMD-DRP detects aggressor rows with a higher precision than SMD-PRP using area-expensive counters. As the 4c-high workloads do not repeatedly activate any single row, SMD-DRP counters rarely exceed the maximum activation threshold, and thus trigger neighbor row refresh only a few times.
7.5.7 SMD’s Performance Sensitivity to Number of Cores
The four-core workloads with high memory intensity (4c-high) largely saturate the DRAM bandwidth offered by four channels. Therefore, four cores are sufficient to capture the performance impact of SMD for cases where DRAM bandwidth utilization is high. To demonstrate this, we evaluate the performance of SMD-FR for a refresh period using highly memory-intensive 8-core (i.e., 8c-high) workloads. The speedup is comparable to the speedup of 4c-high workloads. At 32ms refresh period, SMD-FR provides 8.1% average speedup for 8c-high workloads while it provides 8.7% speedup for 4c-high workloads. SMD provides a slightly lower speedup for 8c-high because of the increased memory request queueing delay, as eight-core workloads produce more memory requests than four-core workloads.
8 Discussion
8.1 Prioritizing Memory Requests
To minimize the impact of rejected activate commands on system performance, an SMD chip could apply a different activate command rejection policy that prioritizes activate commands over maintenance operations. We describe and estimate the performance impacts of one such policy.
“Pause Maintenenace” Policy (SMD-PMP). The key idea of SMD-PMP is to pause an ongoing maintenance operation for a lock region when the memory controller issues an activate command to the lock region. The SMD chip could resume the maintenance operation when the memory controller precharges the bank (i.e., finishes accessing the lock region). For example, the SMD chip sequentially refreshes eight rows in a lock region (section 5.1) during a periodic refresh maintenance operation. If the chip receives an activate command to this lock region while only four out of eight rows have been refreshed, the chip does not continue refreshing the fifth row, but yields control of the lock region to the memory controller.
Latency of Pausing Maintenance Operations. Even if the SMD chip decides to pause the ongoing maintenance operation, the activate command cannot immediately proceed because the lock region might not be in the precharged state (i.e., the DRAM bank might not be ready to activate a row). The DRAM chip requires up to (section 2.1) to bring the under-maintenance lock region’s state to the precharged state (in case an activate command was just issued by the maintenance mechanism), depending on the lock region’s state when the activate command is received.
Estimating the Performance of SMD-PMP. We model SMD-PMP-FR as SMD-FR that refreshes one DRAM row in a lock region before releasing the lock (i.e., SMD-PMP-FR locks a region for at most ). Fig. 18 shows weighted speedup for 60 four-core workloads (20 per memory intensity level) on the left and the number of ACT_NACKs over the number of issued activate commands (which we call “ACT_NACK rate”), averaged across all workloads in every intensity level on the right.
We make two key observations from Fig. 18. First, SMD-PMP-FR provides a larger average speedup than SMD-FR for 4c-high workloads. SMD-PMP-FR’s ACT_NACK rate is 2% smaller than SMD-FR’s, which explains the 0.4% larger average speedup it provides over SMD-FR. Second, SMD-FR-1LR induces a smaller average system slowdown than SMD-PMP-FR-1LR. We conclude that an SMD-PMP design (whose performance could be accurately estimated using our methodology) that pauses maintenance operations to serve activate commands can improve performance.
9 Related Work
This is the first paper to enable DRAM chips that autonomously and efficiently perform various maintenance operations with simple changes to existing DRAM interfaces. No prior work proposes setting the MC free from managing DRAM maintenance operations nor studies the system-level performance and energy impact of autonomous maintenance mechanisms. We briefly discuss relevant prior works.
Changing the DRAM Interface. Several prior works [129, 130, 131, 132] propose using high-speed serial links and packet-based protocols in DRAM chips. SMD differs from prior works in two key aspects. First, none of these works describe how to implement maintenance mechanisms completely within DRAM. We propose five new SMD-based maintenance mechanisms (section 5). Second, prior works significantly overhaul the existing DRAM interface, which makes their proposals more difficult to adopt compared to SMD, which adds only a single ACT_NACK signal to the existing DRAM interface and requires slight modifications in the MC.
Mitigating DRAM Refresh Overhead. Many previous works [16, 17, 15, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43] propose techniques to reduce DRAM refresh overhead. SMD enables maintenance mechanisms for reducing DRAM refresh overhead and other efficient maintenance mechanisms for improving DRAM reliability and security (section 5).
RowHammer Protection. Many prior works [133, 49, 50, 54, 48, 53, 134, 45, 135, 10, 75, 136, 59, 58, 57, 137, 62, 63, 138, 139] propose techniques for RowHammer protection. One can use SMD to implement existing or new RowHammer protection mechanisms. Our goal is not to prove that SMD-DRP is the best RowHammer mitigation mechanism.
Memory Scrubbing. Although prior works report that the overhead of memory scrubbing is small as low scrubbing rate (e.g., scrubbing period of 24 hours [69, 70, 72], 45 minutes per 1 GB scrubbing rate [66], only when idle [67]) is generally sufficient, the cost of scrubbing can dramatically increase for future DRAM chips due to increasing DRAM bit error rate and increasing DRAM chip density. SMD-MS enables efficient scrubbing by eliminating off-chip data transfers.
Leveraging Subarray-level Parallelism. Prior works overlap the latency of accessing multiple subarrays by modifying the DRAM architecture [79, 80, 16]. SMD’s maintenance-access parallelization builds on the basic design proposed in SALP [79] and refresh-access parallelization introduced in [16, 80]. A DDRx interface based on these prior works [79, 80, 16] still needs to change for each new maintenance operation that the DRAM manufacturer needs to implement. SMD, in contrast, allows the implementation of new maintenance mechanisms without making any further changes to the DDRx interface.
Compute Express Link (CXL) [140]. CXL is a cache-coherent interconnect for computing systems. CXL does not define the interface between a memory controller and a DRAM module/chip. Even with CXL, the memory controller chip has to deal with the management complexity of DRAM. SMD can be used in conjunction with CXL to ease management complexity in computing systems.
10 Conclusion
To set the memory controller free from managing DRAM maintenance operations, Self-Managing DRAM (SMD) introduces minimal changes to the existing DRAM chips and memory controllers. SMD enables in-DRAM maintenance operations with no further changes to the DRAM interface, memory controller, or other system components. Using SMD, we implement efficient maintenance mechanisms for DRAM refresh, RowHammer protection, and memory scrubbing. We show that these mechanisms altogether enable a higher performance, more energy efficient and at the same time more reliable and secure DRAM system. We believe and hope that SMD will enable practical adoption of innovative ideas in DRAM design.